OpenStudy (anonymous):

Solve x2 - 4x - 7 = 0 by completing the square

OpenStudy (anonymous):

The main rule is ax2+bx+c right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The general rule for completing the square is: x2+bx=(x+b/2)2-(b/2)

OpenStudy (anonymous):

So it is: (x+4/2)2-(4/2)=7

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Sorry I forgot the square in : -(4/2)

OpenStudy (anonymous):

so it is: (x+2)2-2^2=7

OpenStudy (anonymous):

(x+2)2-4=7

OpenStudy (anonymous):

Take the -4 on to the other side so it becomes: (x+2)2=7+4 (x+2)2=11

OpenStudy (anonymous):

you need to find x so to remove the square u need to multiply both sides by square root

OpenStudy (anonymous):

It becomes: x+2=square root of 11

OpenStudy (anonymous):

move the 2 to the other side and it becomes: x=square root of 11+or- 2

OpenStudy (anonymous):

Because it is a quadratic equation so it must have two solutions

OpenStudy (anonymous):

Hope u understood